# Energy Confinement Time Scaling for the Negative Triangularity Scenario in DIII-D

### v1

@@ Abstract
Abstract — Results from the 2023 negative triangularity campaign on DIII-D demonstrate encouraging energy confinement properties, similar to or exceeding the scaling of the IPB98(y,2) law. This paper describes the procedure with which a new scaling law was regressed specifically from the data from the DIII-D campaign. Given the relatively small size of the single-machine dataset, measures were taken to minimize sampling bias and give a realistic estimate of the large uncertainties from the regression. The resulting power law shows a robustly stronger dependence on plasma current and more severe power degradation as compared to the H-mode scaling law.
@@ I Introduction
The negative triangularity (NT) scenario was further explored in the dedicated multi-week campaign at DIII-D in 2023. This campaign demonstrated the simultaneous attainment of $H_{98y2}>1$, $f_{\mathrm{GW}}>1$, and $\beta_{N}>2.5$ Thome et al. (2024). Of particular note for this paper is the energy confinement performance of NT compared to the IPB98(y,2) scaling law. This scenario showed H-mode-like confinement scaling with the robust avoidance of edge-localized modes, illustrating the potential of NT as a reactor-relevant scenario.
Fig. 1(a) shows the consistency of the dataset’s (thermal) energy confinement time, $\tau_{E,th}$ against the IPB98(y,2) scaling law. It can be seen that NT largely follows H-mode-like scaling, with a divergence at high $\tau_{E,th}$. Additionally, 1(b) shows a similar plot with the measured energy confinement time, $\tau_{E}$, against the prediction of the $L_{89}$ L-mode scaling law. Note that the distinction between $\tau_{E,th}$ and $\tau_{E}$ is that the contribution of the fast particles is subtracted from the former. This figure shows a clear deviation of the confinement time from the L-mode scaling, where higher $\tau_{E}$ approach $\sim 2\times$ the performance as predicted by the $L_{89}$ scaling law. The comparisons in Fig. 1 clearly demonstrate the different behavior of NT confinement time scaling compared to the standard H-mode and L-mode scaling laws, motivating the need for a new engineering scaling regression specifically for NT.
This report provides an extension to the regression results published in Paz-Soldan et al. (2024), giving additional detail as to how the regression and uncertainty quantification were performed — this is covered in the following section. Section III discusses potential future work to further elucidate the NT confinement time scaling law. The scripts used in this analysis are available on GitHub.
@@ II Scaling Law
$\begin{split}\log{\tau_{E,th}}&=\log{C}\\ &+\alpha_{I}\log{I_{p}}+\alpha_{B}\log{B_{t}}+\alpha_{n}\log{n_{e}}+\alpha_{P}\log{P_{\mathrm{tot}}}\\ &+\alpha_{\kappa}\log{\kappa}+\alpha_{\epsilon}\log{\epsilon}+\alpha_{R}\log{R_{0}}+\alpha_{M}\log{M}\end{split}$ (2)
The NT database in the supplementary material of Paz-Soldan et al. (2024) is utilized, with additional filters applied. Specifically, stationary phases with too high $\dot{W}$, too high saturated $n=2$ mode activity, $f_{GW}>1$, and zero torque are omitted. This results in 260 stationary phases to fit to, across 132 discharges. Given all the data in the dataset is from DIII-D with minor changes in the shaping parameters ($\kappa$, $\epsilon$, $R_{0}$), it is not possible to fit to these parameters. Additionally, there was no variation in the mass of the main ion species, $M$, in the whole dataset. As a result, the exponents for these parameters are fixed to those from the IPB98(y,2) scaling law. The fitting data is displayed in Fig. 2, reproduced from Appendix A of Paz-Soldan et al. (2024).
Figure 2: Data scatter matrix for all data used in the regression. The Pearson correlation coefficient is indicated in the corresponding subplots, and does not show significant co-linearity.
It can be seen in 2 that a uniform distribution does not exist across all the parameters, which would lead to linear least squares having sampling bias. To combat this, kernel density estimation (KDE) is used to assign weights to the data points such that less sampled regions of the space are weighted more strongly and more sampled regions of the space are weighted more weakly. This technique has been used in error field penetration scaling studies to alleviate sampling bias Logan et al. (2020). The regression to Eq. 2 can be seen in Fig. 3.
@@
A simple yet effective method for quantifying the uncertainties in this regression is by bootstrapping, which involves performing numerous regressions on a randomly sampled subset of the full dataset. Fig. 4 illustrates this technique on 1000 regressions of 75% and 50% randomly sampled subsets. It is clear that there is a large spread in the values of the exponents given the choice of data, but clear trends can be seen in the exponents compared to those from IPB98(y,2). Namely, there is a stronger dependence on $I_{p}$ and $P_{\mathrm{tot}}$. To capture the largest reasonable uncertainties, we tentatively take the mean of each quantity’s 50% resampled distribution plus/minus $2\sigma$, obtaining a 95% confidence interval. The results of this are summarized in Table 1.
@@ III Future Work
Given that the regression results presented in the previous section were only using DIII-D data, scaling with the shaping parameters ($\kappa$, $\epsilon$, and $R_{0}$) was not possible due to their low variance. As such, a natural extension would be to include NT data from additional machines that expands the dataset in the operating space of these shaping parameters. This may introduce co-linearities given the correlation between $\epsilon$ and $R_{0}$, which could be addressed using the ridge regression technique.
This work has exclusively performed weighted linear least squares regression, which while straightforward, may not be the robust technique to use. Indeed, the work in Verdoolaege et al. (2021) has taken extensive effort to robustly regress a large multi-machine dataset using Bayesian techniques which properly addresses sampling bias, co-linearities, and uncertainty quantification. Implementing this technique for NT scaling would be a natural next step.

### v2

# Energy Confinement Time Scaling Law Derived from Paz-Soldan NF 2024

@@ Abstract
Abstract — Results from C. Paz-Soldan et al 2024 Nucl. Fusion 64 094002 Paz-Soldan et al. (2024) demonstrate encouraging energy confinement properties for the negative triangularity scenario, similar to or exceeding the scaling of the IPB98(y,2) law. This work describes the procedure with which a new scaling law was regressed specifically from the data from the aforementioned article. Given the relatively small size of the single-machine dataset, kernel density estimation was employed to minimize sampling bias and bootstrapping was used to give a realistic estimate of the large uncertainties from the regression. The resulting power law shows a robustly stronger dependence on plasma current and more severe power degradation as compared to the H-mode scaling law.
@@ I Introduction
The negative triangularity (NT) scenario was further explored in dedicated studies in Paz-Soldan et al. (2024); Thome et al. (2024). These studies demonstrated the simultaneous attainment of $H_{98y2}>1$, $f_{\mathrm{GW}}>1$, and $\beta_{N}>2.5$. Of particular note for this work is the energy confinement performance of NT compared to the IPB98(y,2) scaling law ITER Physics Expert Group on Confinement and Transport (1999). The NT scenario showed H-mode-like confinement scaling with the robust avoidance of edge-localized modes, illustrating the potential of NT as a reactor-relevant scenario.
The NT database in the supplementary material of Ref. Paz-Soldan et al. (2024) is utilized. Fig. 1(a) shows the consistency of the dataset’s (thermal) energy confinement time, $\tau_{E,th}$, against the IPB98(y,2) scaling law. It can be seen that NT largely follows H-mode-like scaling, with a divergence at high $\tau_{E,th}$. Additionally, Fig. 1(b) shows a similar plot with the measured energy confinement time, $\tau_{E}$, against the prediction of the $L_{89}$ L-mode scaling law. Note that the distinction between $\tau_{E,th}$ and $\tau_{E}$ is that the contribution of the fast particles is subtracted from the former. This figure shows a clear deviation of the confinement time from the L-mode scaling, where higher $\tau_{E}$ approach $\sim 2\times$ the performance as predicted by the $L_{89}$ scaling law. The comparisons in Fig. 1 clearly demonstrate the different behavior of NT confinement time scaling compared to the standard H-mode and L-mode scaling laws, motivating the need for a new engineering scaling regression specifically for NT.
This report provides an extension to the regression results published in Ref. Paz-Soldan et al. (2024), giving additional detail on the regression procedure and uncertainty quantification — this is covered in the following section. Section III discusses potential future work to further elucidate the NT confinement time scaling law. The scripts used in this analysis are available on GitHub.
@@ II Scaling Law
$\begin{split}\log{\tau_{E,th}}&=\log{C}\\ &+\alpha_{I}\log{I_{p}}+\alpha_{B}\log{B_{t}}+\alpha_{n}\log{n_{e}}+\alpha_{P}\log{P_{\mathrm{tot}}}\\ &+\alpha_{\kappa}\log{\kappa}+\alpha_{\epsilon}\log{\epsilon}+\alpha_{R}\log{R_{0}}+\alpha_{M}\log{M}.\end{split}$ (2)
As previously mentioned, the NT database in the supplementary material of Ref. Paz-Soldan et al. (2024) is utilized, with additional filters applied. Specifically, stationary phases with too high $\dot{W}$, too high saturated $n=2$ mode activity, $f_{GW}>1$, and zero torque are omitted. This results in 260 stationary phases to fit to, across 132 discharges. Given all the data in the dataset is from a single device with minor changes in the shaping parameters ($\kappa$, $\epsilon$, $R_{0}$), it is not possible to fit to these parameters. Additionally, there was no variation in the mass of the main ion species, $M$, in the whole dataset. As a result, the exponents for these parameters are fixed to those from the IPB98(y,2) scaling law. The fitting data is displayed in Fig. 2, reproduced from Appendix A of Ref. Paz-Soldan et al. (2024).
Figure 2: Data scatter matrix for all data used in the regression, reproduced from Appendix A of Ref. Paz-Soldan et al. (2024). The Pearson correlation coefficient is indicated in the corresponding subplots, and does not show significant co-linearity.
It can be seen in Fig. 2 that a uniform distribution does not exist across all the parameters, which would lead to linear least squares having sampling bias. To combat this, kernel density estimation (KDE) is used to assign weights to the data points such that less sampled regions of the space are weighted more strongly and more sampled regions of the space are weighted more weakly. This technique has been used in error field penetration scaling studies to alleviate sampling bias Logan et al. (2020). The regression to Eq. 2 can be seen in Fig. 3.
@@
A simple yet effective method for quantifying the uncertainties in this regression is by bootstrapping, which involves performing numerous regressions on a randomly sampled subset of the full dataset. Fig. 4 illustrates this technique on 1000 regressions of 75% and 50% randomly sampled subsets. It is clear that there is a large spread in the values of the exponents given the choice of data, but clear trends can be seen in the exponents compared to those from IPB98(y,2). Namely, there is a stronger dependence on $I_{p}$ and $P_{\mathrm{tot}}$. To capture the largest reasonable uncertainties, we tentatively take the mean of each quantity’s 50% resampled distribution plus/minus $2\sigma$, obtaining a 95% confidence interval. The results of the regression and uncertainty quantification are summarized in Table 1.
@@ III Future Work
Given that the regression results presented in the previous section were only using data from a single device, scaling with the shaping parameters ($\kappa$, $\epsilon$, and $R_{0}$) was not possible due to their low variance Paz-Soldan et al. (2024). As such, a natural extension would be to include NT data from additional machines that expands the dataset in the operating space of these shaping parameters. This may introduce co-linearities given the correlation between $\epsilon$ and $R_{0}$, which could be addressed using the ridge regression technique.
This work has exclusively performed weighted linear least squares regression, which, while straightforward, may not be the robust technique to use. Indeed, the work in Ref. Verdoolaege et al. (2021) has taken extensive effort to robustly regress a large multi-machine dataset using Bayesian techniques which properly addresses sampling bias, co-linearities, and uncertainty quantification. Implementing this technique for NT scaling would be a natural next step.